# Mitigating distribution shift in machine learning-augmented hybrid simulation

### v1

@@ 3.1 Distribution shift in linear dynamics and motivation of tangent-space regularized estimator
As there exists few tools to tackle the general non-linear dynamics Eq. 1, we switch to the simpler case of linear dynamics, which also capture some key features in general situations. We consider the following hybrid simulation problem
$\partial_{t}\mathbf{u}$ $=A\mathbf{u}+B\mathbf{y},\quad\mathbf{u}\in\mathbb{R}^{m},\mathbf{y}\in\mathbb{R}^{n},A\in\mathbb{R}^{m\times m},B\in\mathbb{R}^{m\times n}$ (12)
@@
We assume that the state variables $\mathbf{u}_{i}$s in training data come from several trajectories and $\mathbf{y}_{i}$s may be subject to some random measurement error, i.e. $\mathbf{y}_{i}=C^{*}\mathbf{u}_{i}+\epsilon_{i}$, or in matrix form $Y=C^{*}\mathbf{U}+\epsilon$. The key point is that these trajectories may belong to a low-dimensional subspace in the high-dimensional state space. This is the case for many scientific computing problems, such as computational fluid dynamics [38], and is also known more generally as the manifold hypothesis [34]. We use $V\subset\mathbb{R}^{m}$ to denote the subspace that contains all the training data $\mathbf{u}_{i}$. In this setting, the appearance of this low-dimensional structure may be caused by two different situations. In the first situation, the initial value of $\mathbf{u}$ is supported on the subspace spanned by several eigenvectors of the evolution operator $e^{(A+BC^{*})}$, i.e.
$\mathbf{u}_{0}\in\operatorname{span}\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l}\},\quad e^{(A+BC^{*})}\mathbf{v}_{1}=\lambda_{1}\mathbf{v}_{1},\cdots,e^{(A+BC^{*})}\mathbf{v}_{l}=\lambda_{l}\mathbf{v}_{l},$ (13)
Then, all the training data $\mathbf{u}_{i}$ belong to this subspace, i.e. $V=\operatorname{span}\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l}\}$. The second situation is that the dynamics is degenerate in the sense that the evolution matrix $A$ is not of full rank. This will not appear in the differential formulation as infinitesimal transformations in the form $e^{(A+BC^{*})\Delta t}$ are always non-degenerate. However, for a discrete dynamics
$\mathbf{u}_{k+1}=A\mathbf{u}_{k}+B\mathbf{y}_{k},\quad\mathbf{y}_{k}=C^{*}\mathbf{u}_{k},$ (14)
if the matrix $A+BC^{*}$ is degenerate, then all the data $\mathbf{u}_{k}$ will belong to the range of this operator $A+BC^{*}$, which is a proper subspace. Most cases of scientific computing applications belong to the first class. Due to the degeneracy in the training data, the least squares solution is not unique. Hence, various empirical regularizations (e.g. $\ell^{2}$ regularization) can be introduced to obtain an estimator with desirable properties.
@@
$l_{OLS}(\widehat{C})=\mathbb{E}\left\|(\widehat{C}-C^{*})\mathbf{u}\right\|^{2},$ (15)
where $\mathbf{u}$ is sampled from the low dimensional subspace $V$ and follows the same distribution as the training data. We refer to this as statistical estimation error. However, we are actually interested in the error of the simulated dynamics
$\partial_{t}\widehat{\mathbf{u}}=A\widehat{\mathbf{u}}+B\widehat{\mathbf{y}},\quad\widehat{\mathbf{y}}=\widehat{C}\widehat{\mathbf{u}},$ (16)
@@
$\partial_{t}(\widehat{\mathbf{u}}-\mathbf{u})=$ $\ (A+B\widehat{C})\widehat{\mathbf{u}}-(A+BC^{*})\mathbf{u}$ (17)
$=$ $\ A(\widehat{\mathbf{u}}-\mathbf{u})+B(\widehat{C}\widehat{\mathbf{u}}-\widehat{C}P_{V}\widehat{\mathbf{u}}+\widehat{C}P_{V}\widehat{\mathbf{u}}-\widehat{C}\mathbf{u}+\widehat{C}\mathbf{u}-C^{*}\mathbf{u})$
$=$ $\ (A+B\widehat{C}P_{V})(\widehat{\mathbf{u}}-\mathbf{u})+B(\widehat{C}-C^{*})\mathbf{u}+B\widehat{C}(\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}}),$
@@
where $P_{V}$ is the orthogonal projection onto the data subspace $V$ and we have $P_{V}\mathbf{u}=\mathbf{u}$ since $\mathbf{u}$ belongs to the data subspace. As derived in Eq. 17, the overall error of the trajectory during evolution can be decomposed into three parts. In the following we make a detailed analysis of each term.
The first term can be viewed as an amplitude factor of the error propagation and can be easily bounded by
$\left\|(A+B\widehat{C}P_{V})(\widehat{\mathbf{u}}-\mathbf{u})\right\|_{2}\leq\left\|A+B\widehat{C}P_{V}\right\|_{2}\left\|\widehat{\mathbf{u}}-\mathbf{u}\right\|_{2}.$ (18)
The operator $A+B\widehat{C}P_{V}$ can be viewed as a posterior estimation of the error propagation and a stability constant of the simulated dynamics.
@@
The second term is simply the statistical estimation error of the estimator $\widehat{C}$ defined in Eq. 15. Given that the initial condition $\mathbf{u}_{0}$ of the training and test trajectories are sampled from the same distribution, this term is well-bounded in classical statistical learning theory [20], provided the number of data tuples is sufficiently large. In particular, this error has nothing to do with distribution shift.
@@
Now we move to the last part $B\widehat{C}(\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}})$, which can be more easily understood from a geometric viewpoint. The last factor $(\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}})$ measures how far the simulated trajectory is away from the data subspace $V$. In other words, this term measures how far the simulated trajectory’s state distribution “shifts” from the true distribution. This is exactly the term we want to obtain to control the distribution shift in simulation. Meanwhile, this term forms sharp contrast with previous two terms in the sense that it is not automatically bounded in most algorithm for estimating $\widehat{C}$. We will illustrate this point further by deriving a error bound of the simulated dynamics, which shows that it depends sensitively on this distribution shift term. Thus, we can improve prediction accuracy if we combat such distribution shifts.
Now, taking the expectation of the inner product between $\widehat{\mathbf{u}}-\mathbf{u}$ and Eq. 17 and applying Gronwall’s lemma one obtains the following bound on the error propagation.
$\mathbb{E}\left\|\widehat{\mathbf{u}}(T)-\mathbf{u}(T)\right\|_{2}^{2}\leq$ (19)
$\frac{\left\|B\right\|_{2}^{2}\mathbb{E}\left\|(\widehat{C}-C^{*})\mathbf{u}\right\|^{2}+\left\|B\widehat{C}\right\|_{2}^{2}\sup_{t\leq T}\left\|\widehat{\mathbf{u}}(t)-P_{V}\widehat{\mathbf{u}}(t)\right\|_{2}^{2}}{\left\|A+B\widehat{C}P_{V}\right\|_{2}^{2}+3}(e^{(\left\|A+B\widehat{C}P_{V}\right\|_{2}^{2}+3)T}-1).$
@@
An interpretation for this result is that the trajectory error can be divided into two parts: $\left\|B\right\|_{2}^{2}\mathbb{E}\left\|(\widehat{C}-C^{*})\mathbf{u}\right\|^{2}$ and $\left\|B\widehat{C}\right\|_{2}^{2}\sup_{t}\left\|\widehat{\mathbf{u}}(t)-P_{V}\widehat{\mathbf{u}}(t)\right\|_{2}^{2}$. The first error is the statistical estimation error of unresolved component, which can be reduced by better training algorithms, better models, or more data etc. This is the familiar term one faced in supervised learning under common i.i.d. setting. The second error relates to the distribution shift since it is measured by the difference between the sampled trajectories and underlying data manifold, i.e. $\left\|\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}}\right\|=\operatorname{dist}(\widehat{\mathbf{u}},V)$. Moreover, this shift is caused by the interaction between the estimator of unresolved model and the resolved model dynamics. Specifically, the estimator makes some error in prediction and this error is propagated via the resolved model dynamics and then fed into the next round of hybrid simulation. This distribution shift in MLHS is far different from the one in computer science such as covariate shift and label shift. In computer vision, the cause of distribution shift may be extrinsic [23], i.e. the pictures in training set are all taken at daytime while those for testing are all taken at night. Such shifts are hard to model, so their resolution tends to depend on data augmentation and related techniques, e.g. the Dagger algorithm [41]. Instead of changing the optimization problem, the data source is modified. This shares much similarity with adversarial training, which also add more data to the training set to make the prediction robust under adversarial attack, a typical distribution shift in the area of computer vision [15]. In contrast, in our setting the distribution shift is intrinsically driven by the hybrid simulation structure of which we have partial knowledge. Hence, we can quantify this distribution shift and design specific algorithm to mitigate them.
@@
Returning to the linear problem, in order to guarantee that the error $\widehat{\mathbf{u}}-\mathbf{u}$ is bounded over the simulated trajectory, a natural choice is to use a regularization for $\left\|\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}}\right\|_{2}$. The most naive choice would be $\left\|\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}}\right\|_{2}$ itself, but a problem is that $\widehat{\mathbf{u}}$ is calculated via hybrid simulation until time $t$, which corresponds to a rather complicated computational graph. This makes subsequent gradient-based optimization computational expensive. Hence, we take the advantage of an one-step predictor, i.e. we set
$\widehat{\mathbf{u}}_{k+1}=\mathbf{u}_{k}+\Delta t(A\mathbf{u}_{k}+B\widehat{C}\mathbf{u}_{k}).$ (20)
The $\widehat{\mathbf{u}}_{k+1}$ is calculated based on a single step of the simulated dynamics from the ground truth solution $\mathbf{u}_{k}$. Assume $\mathbf{u}_{k}\in V$, one has
$\left\|\widehat{\mathbf{u}}_{k+1}-P_{V}\widehat{\mathbf{u}}_{k+1}\right\|_{2}=\Delta t\left\|P_{V^{\perp}}(A\mathbf{u}_{k}+B\widehat{C}\mathbf{u}_{k})\right\|_{2},$ (21)
@@
and we may penalize the right hand side to promote trajectory alignment with the data manifold. Wrapping up all the ingredients, we state the loss function in the linear case as
$\min_{C}l(C):=\min_{C}\mathbb{E}_{(\mathbf{u},\mathbf{y})}\left(\left\|\mathbf{y}-C\mathbf{u}\right\|_{2}^{2}+\lambda\left\|P_{V^{\perp}}(A+BC)\mathbf{u}\right\|_{2}^{2}\right),$ (22)
@@
and here $\lambda$ is a penalty strength parameter which can be chosen during training. We briefly discuss in the supplementary materials LABEL:lambda how to choose $\lambda$.
@@
In above discussion, we have assumed that the underlying data subspace $V$ is known a priori or we can calculate $P_{V^{\perp}}$ directly. However, in most cases the only information we have is a set of resolved state variables $\mathbf{u}_{i}$s. Then, we can use standard dimension reduction methods such as principle component analysis [20] to obtain a subspace $V$, which can be thought as an approximation of the data manifold. The loss objective Eq. 22 can thus be calculated.
@@ 3.2 Algorithm in the general form
The discussion above applies to linear dynamics where the underlying data manifold is also linear. In this subsection we generalize it to non-linear cases and state the general form of our algorithm.
@@
Let us return to our original formulation of the task Eq. 1, where we have a dataset $\left\{(\mathbf{u}_{1},\mathbf{y}_{1},t_{1}),(\mathbf{u}_{2},\mathbf{y}_{2},t_{2}),\cdots,(\mathbf{u}_{N},\mathbf{y}_{N},t_{N})\right\}$. As before, we assume that the resolved variable lies on a low-dimensional manifold $\mathcal{M}$ in the configuration space and the estimator $\phi_{\theta}(\mathbf{u},t)$ achieves high accuracy along this whole manifold while we have no guarantee outside the manifold. The key difference between the general case and the linear case is that this manifold may not be linear and we do not have an orthogonal decomposition.
@@
To resolve this, we leverage an autoencoder to represent the data manifold structure, i.e. we learn two neural networks $E:\mathbb{R}^{m}\rightarrow\mathbb{R}^{l}$ (encoder) and $D:\mathbb{R}^{l}\rightarrow\mathbb{R}^{m}$ (decoder) to represent the manifold. The encoder maps the state variable $\mathbf{u}$ into a low dimensional space and decoder maps them back to original space. We optimize $E,D$ simultaneously to minimize the reconstruction error
$\min_{E,D}\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{u}_{i}-D(E(\mathbf{u}_{i}))\right\|_{2}^{2}.$ (23)
Here, we only use the resolved states $\mathbf{u}_{i}$ from the dataset. After training the autoencoder, we obtain a function
$F:\mathbb{R}^{m}\rightarrow\mathbb{R}_{\geq 0},\quad F(\mathbf{u})=\left\|\mathbf{u}-D(E(\mathbf{u}))\right\|_{2}^{2},$ (24)
which implicitly parametrizes the data manifold via $F(\mathbf{u})=0,\forall\mathbf{u}\in\mathcal{M}$. Notice that the linear version of autoencoder is exactly the principle component analysis we mentioned to identify linear subspace $V$. Such techniques to identify data manifolds have been used in improving robustness against adversarial attacks [5, 6]. Besides parametrizing the data manifold, $F$ can be used as an indicator of distribution shift, i.e. for any state variable $\mathbf{u}$, we use $F(\mathbf{u})=\left\|\mathbf{u}-D(E(\mathbf{u}))\right\|_{2}^{2}$ to measure its distribution shift w.r.t. the training data. This is used in all the figures in this paper to quantify distribution shift. Strictly speaking, training only ensures that the learned manifold is contained in the zero level-set of $F$. Thus, the gradient of $F$ belongs to the normal space of the manifold $\mathcal{M}$ at the point $\mathbf{u}$, i.e. $\nabla F\in\mathcal{M}^{\perp}$. Minimizing the inner product with $\nabla F$ now means partially minimizing the orthogonal components, which is the cause of the distribution shift in our formulation. This strictly generalizes the linear case Eq. 22, where the data subspace $V$ is given by $\operatorname{span}\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l}\}$. Augmenting this basis to obtain an orthonormal basis $\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l},\mathbf{v}_{l+1},\cdots,\mathbf{v}_{m}\}$ for the whole space, a choice of the function $F$ is given by
$F(\mathbf{u})=\frac{1}{2}\sum_{i=l+1}^{m}(\mathbf{u}^{T}\mathbf{v}_{i})^{2},$ (25)
and the gradient of $F$ is
$\nabla F(\mathbf{u})=\left(\sum_{i=l+1}^{m}\mathbf{v}_{i}\mathbf{v}_{i}^{T}\right)\mathbf{u},$ (26)
which is just the orthogonal projection onto $V^{\perp}=\operatorname{span}\{\mathbf{v}_{l+1},\mathbf{v}_{l+2},\cdots,\mathbf{v}_{m}\}.$ We can state the regularized loss function of our algorithm in general case:
$\min_{\theta}l(\theta):=\min_{\phi_{\theta}}\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\mathbf{y}_{i}-\phi_{\theta}(\mathbf{u}_{i})\right\|_{2}^{2}+\lambda\left(\left(\nabla F(\mathbf{u}_{i})\right)^{T}\mathcal{L}(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})\right)^{2}\right).$ (27)
@@
The first term above is the least squares term and the second term measures how far the estimated tangent vector $\partial_{t}\mathbf{u}_{i}=\mathcal{L}(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})$ deviates from the tangent space of $\mathcal{M}$ at $\mathbf{u}_{i}$. In the second term, $F(\mathbf{u})=\left\|\mathbf{u}-D(E(\mathbf{u}))\right\|_{2}^{2}$ and $D,E$ are optimized via Eq. 23 and frozen during the training of $\phi_{\theta}$. Since the loss function is regularized by a term which approximately measures the deviation of the tangent vector $\mathcal{L}(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})$ from the tangent space of the data manifold, we name our algorithm “MLHS with tangent-space regularized estimator”. The overall algorithm can be summarized in Algorithm 1.

### v2

@@ 3.1 Distribution shift in linear dynamics and motivation of tangent-space regularized estimator
As there exist few tools to tackle the general non-linear dynamics Eq. 1, we switch to the simpler case of linear dynamics, which also capture some key features in general situations. We consider the following the hybrid simulation problem
(12) $\frac{d\mathbf{u}}{dt}$ $=A\mathbf{u}+B\mathbf{y},\quad\mathbf{u}\in\mathbb{R}^{m},\mathbf{y}\in\mathbb{R}^{n},A\in\mathbb{R}^{m\times m},B\in\mathbb{R}^{m\times n}$
@@
We assume that the state variables $\mathbf{u}_{i}$s in training data come from several trajectories and $\mathbf{y}_{i}$s may be subject to some random measurement error, i.e. $\mathbf{y}_{i}=C^{*}\mathbf{u}_{i}+\epsilon_{i}$, or in matrix form $Y=C^{*}\mathbf{U}+\epsilon$. The key point is that these trajectories may belong to a low-dimensional subspace in the high-dimensional state space. This is the case for many scientific computing problems, e.g. parametric elliptic PDE [3] and is also related to the notions of Kolmogorov N-width and the manifold hypothesis  [11]. Although transport-dominated problems with low-decaying rates of the Kolmogorov N-width are difficult to approximate using linear reduced-order models, a large amount of literature [20, 21] show evidence of the existence of coherent structures in turbulence and their low-dimensional nature. Moreover, recent work [9] using nonlinear model reduction has achieved efficient reduction over this type of problems.
We use $V\subset\mathbb{R}^{m}$ to denote the subspace that contains all the training data $\mathbf{u}_{i}$. In this setting, the appearance of this low-dimensional structure may be caused by two different situations. In the first situation, the initial value of $\mathbf{u}$ is supported on the subspace spanned by several eigenvectors of the evolution operator $e^{(A+BC^{*})}$, i.e.
(13) $\mathbf{u}_{0}\in\operatorname{span}\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l}\},\quad e^{(A+BC^{*})}\mathbf{v}_{1}=\mathrm{eig}_{1}\mathbf{v}_{1},\cdots,e^{(A+BC^{*})}\mathbf{v}_{l}=\mathrm{eig}_{l}\mathbf{v}_{l},$
Then, all the training data $\mathbf{u}_{i}$ belong to this subspace, i.e. $V=\operatorname{span}\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l}\}$. The second situation is that the dynamics is degenerate in the sense that the evolution matrix $A$ is not of full rank. This will not appear in the differential formulation as infinitesimal transformations in the form $e^{(A+BC^{*})\Delta t}$ are always non-degenerate. However, for a discrete dynamic $\mathbf{u}_{k+1}=A\mathbf{u}_{k}+B\mathbf{y}_{k},\quad\mathbf{y}_{k}=C^{*}\mathbf{u}_{k}$, if the matrix $A+BC^{*}$ is degenerate, then all the data $\mathbf{u}_{k}$ will belong to the range of this operator $A+BC^{*}$, which is a proper subspace. Most cases of scientific computing applications belong to the first class. Due to the degeneracy in the training data, the least squares solution is not unique. Hence, various empirical regularizations (e.g. $\ell^{2}$ regularization) can be introduced to obtain an estimator with desirable properties.
@@
(14) $l_{\mathrm{OLS}}(\widehat{C})=\mathbb{E}\left\|(\widehat{C}-C^{*})\mathbf{u}\right\|_{2}^{2},$
where $\mathbf{u}$ is sampled from the low dimensional subspace $V$ and follows the same distribution as the training data. We refer to this as the statistical estimation error or a priori error. However, we are interested in the a posteriori error of the simulated dynamics
(15) $\frac{d\widehat{\mathbf{u}}}{dt}=A\widehat{\mathbf{u}}+B\widehat{\mathbf{y}},\quad\widehat{\mathbf{y}}=\widehat{C}\widehat{\mathbf{u}},$
@@
(16) $\frac{d}{dt}(\widehat{\mathbf{u}}-\mathbf{u})=$ $\ (A+B\widehat{C})\widehat{\mathbf{u}}-(A+BC^{*})\mathbf{u}$
$=$ $\ (A+B\widehat{C})\operatorname{P}_{V}(\widehat{\mathbf{u}}-\mathbf{u})+B(\widehat{C}-C^{*})\mathbf{u}+(A+B\widehat{C})(\widehat{\mathbf{u}}-\operatorname{P}_{V}\widehat{\mathbf{u}}),$
@@
where $\operatorname{P}_{V}$ is the orthogonal projection onto the data subspace $V$, and we have $\operatorname{P}_{V}\mathbf{u}=\mathbf{u}$ since $\mathbf{u}$ belongs to the data subspace. As derived in Eq. 16, the overall error of the trajectory during evolution can be decomposed into three parts. In the following, we make a detailed analysis of each term.
The first term can be viewed as an amplitude and stability factor of the error propagation of the simulated dynamics inside the data manifold $V$. If we assume the ground truth dynamics is stable in $V$, while not necessary to be stable over the whole state space, then this term can be simply bounded by the spectral information of the generator $(A+B\widehat{C})\operatorname{P}_{V}$, see details in 3.5.
@@
The second term is simply the statistical estimation error of the estimator $\widehat{C}$ defined in Eq. 14. Given that the initial condition $\mathbf{u}_{0}$ of the training and test trajectories are sampled from the same distribution, this term is well-bounded in classical statistical learning theory [19], provided the number of data tuples is sufficiently large. In particular, this error has nothing to do with the distribution shift.
@@
Now we move to the last part $(A+B\widehat{C})(\widehat{\mathbf{u}}-\operatorname{P}_{V}\widehat{\mathbf{u}})$, which can be more easily understood from a geometric viewpoint. The last factor $(\widehat{\mathbf{u}}-\operatorname{P}_{V}\widehat{\mathbf{u}})$ measures how far the simulated trajectory is away from the data subspace $V$, i.e. $\left\|\widehat{\mathbf{u}}-\operatorname{P}_{V}\widehat{\mathbf{u}}\right\|_{2}=\operatorname{dist}(\widehat{\mathbf{u}},V)$. In other words, this term measures how far the simulated trajectory’s distribution “shifts” from the true distribution. This is exactly the term we want to control to maintain simulation fidelity. This term forms a sharp contrast with the two terms previously mentioned in the sense that it is not automatically bounded in most algorithms for estimating $\widehat{C}$. We will illustrate this point further by deriving an error bound of the simulated dynamics, showing its sensitive dependency on the distribution shift term. Thus, we can improve a posteriori accuracy if we combat such distribution shifts.
Now, we can formally write the error propagation along the simulated dynamics as
(17) $\ \widehat{\mathbf{u}}(T)-\mathbf{u}(T)$
[…]
@@
The exponential factor $e^{(A+B\widehat{C})\operatorname{P}_{V}(T-t)}$ corresponds to the first term while two terms in the bracket that follows correspond to the second and the third terms mentioned above. The distribution shift term in MLHS, i.e. the second term in the bracket is far different from classical machine learning concepts such as covariate shift and label shift. In computer vision, the cause of distribution shift may be extrinsic [25], i.e. the pictures in training sets are all taken during the daytime while those for testing are all taken at night. Such shifts are hard to model, so their resolution tends to depend on data augmentation and related techniques, e.g. the Dagger algorithm [23, 47], where instead of changing the optimization problem, the data source is modified. This shares many similarities with adversarial training, which also adds more data to the training set to make the prediction robust under adversarial attack, a typical distribution shift in the area of computer vision [15]. In contrast, in our setting, the distribution shift is intrinsically driven by the hybrid simulation structure of which we have partial knowledge. Hence, we can quantify this distribution shift and design specific algorithms to mitigate them.
@@
Returning to the linear problem, in order to guarantee that the error $\widehat{\mathbf{u}}-\mathbf{u}$ is bounded over the simulated trajectory, a natural choice is to use a regularization for $\left\|\widehat{\mathbf{u}}-\operatorname{P}_{V}\widehat{\mathbf{u}}\right\|_{2}$. The most naive choice would be $\left\|\widehat{\mathbf{u}}-\operatorname{P}_{V}\widehat{\mathbf{u}}\right\|_{2}$ itself, but a problem is that $\widehat{\mathbf{u}}$ is calculated via hybrid simulation until time $t$, which corresponds to a rather complicated computational graph. This makes subsequent gradient-based optimization computationally expensive. Hence, we take advantage of a one-step predictor, where we set $\widehat{\mathbf{u}}_{k+1}=\mathbf{u}_{k}+(A\mathbf{u}_{k}+B\widehat{C}\mathbf{u}_{k})\Delta t$, meaning that $\widehat{\mathbf{u}}_{k+1}$ is calculated based on a single step of the simulated dynamics from the ground truth solution $\mathbf{u}_{k}$. Assuming $\mathbf{u}_{k}\in V$, one has
(18) $\left\|\widehat{\mathbf{u}}_{k+1}-\operatorname{P}_{V}\widehat{\mathbf{u}}_{k+1}\right\|_{2}=\Delta t\left\|\operatorname{P}_{V^{\perp}}(A\mathbf{u}_{k}+B\widehat{C}\mathbf{u}_{k})\right\|_{2},$
@@
and we may penalize the right-hand side to promote trajectory alignment with the data subspace. Wrapping up all the ingredients, we state the loss function in the linear case as
(19) $\min_{C}l(C):=\min_{C}\mathbb{E}_{(\mathbf{u},\mathbf{y})}\left(\left\|\mathbf{y}-C\mathbf{u}\right\|_{2}^{2}+\lambda\left\|\operatorname{P}_{V^{\perp}}(A+BC)\mathbf{u}\right\|_{2}^{2}\right),$
@@
and here $\lambda$ is a penalty strength parameter that can be chosen during training. We briefly discuss in subsection 2.5 of supplementary materials on choosing $\lambda$.
@@
In the above discussion, we have assumed that the underlying data subspace $V$ is known a priori, or we can calculate $\operatorname{P}_{V^{\perp}}$ directly. However, in most cases, the only information we have is a set of resolved state variables $\mathbf{u}_{i}$s. Then, we can use standard dimension reduction methods such as principal component analysis [19] to obtain a subspace $V$, which can be thought of as an approximation of the data subspace. The loss objective Eq. 19 can thus be calculated.
@@ 3.2 Algorithm in the general form
The discussion above applies to linear dynamics where the data also lies in a linear subspace. In this subsection, we generalize it to non-linear cases and state the general form of our algorithm.
@@
Let us return to our original formulation of the task Eq. 1, where we have a dataset $\left\{(\mathbf{u}_{1},\mathbf{y}_{1},t_{1}),(\mathbf{u}_{2},\mathbf{y}_{2},t_{2}),\cdots,(\mathbf{u}_{N},\mathbf{y}_{N},t_{N})\right\}$. As before, we assume that the resolved variable $\mathbf{u}$ lies on a $l$-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{m}$ in the configuration space and the estimator $\phi_{\theta}(\mathbf{u},t)$ achieves high accuracy along this manifold while it has no guarantee outside the manifold. The key difference between the general and the linear dynamics is that $\mathcal{M}$ may not be linear and we do not have a linear projection onto it.
@@
To resolve this, we leverage an autoencoder to represent the data manifold structure: we learn two neural networks $E:\mathbb{R}^{m}\rightarrow\mathbb{R}^{l}$ (encoder) and $D:\mathbb{R}^{l}\rightarrow\mathbb{R}^{m}$ (decoder) to represent the manifold. $E,D$ are simultaneously optimized to minimize the reconstruction error
(20) $\min_{E,D}\frac{1}{N}\sum_{i=1}^{N}F^{2}(\mathbf{u}_{i}),\quad F(\mathbf{u}_{i}):=\left\|\mathbf{u}_{i}-D(E(\mathbf{u}_{i}))\right\|_{2}.$
Omitting optimization issues, we can assume that after the training, the function $F$ vanishes on $\mathcal{M}$ and a greater value of $F$ implies a greater deviation to the manifold. Thus $F$ can be used as an indicator of distribution shift and is used in all the figures in this paper to quantify it.
The nonlinear analog of the regularization term in Eq. 19 should be the projection to the normal space $N_{\mathbf{u}_{i}}\mathcal{M}=T_{\mathbf{u}_{i}}^{\perp}\mathcal{M}$ at $\mathbf{u}_{i}$. However, constructing this projection operator would require a highly accurate optimized decoder map [26] so that $dD_{\mathbf{z}_{i}}(\mathbf{e}_{i}),i=1,...,l$ forms a basis of the tangent space $T_{\mathbf{u}_{i}}\mathcal{M}$. Here $dD_{\mathbf{u}_{i}}$ is the differential of the mapping $D$ at $\mathbf{z}_{i}$ and $\{\mathbf{e}_{i},i=1,...l\}$ is the standard basis of $\mathbb{R}^{l}$. To reduce the computational cost, we pick the normal vector $\nabla F(\mathbf{u}_{i})\in N_{\mathbf{u}_{i}}\mathcal{M}$ and use $(\nabla F(\mathbf{u}_{i}))^{T}L(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})$ to measure the deviation of the velocity along this normal direction. The fact that $\nabla F(\mathbf{u}_{i})$ belongs to the normal space of $\mathcal{M}$ at $\mathbf{u}_{i}$ directly follows from the assumption that $F$ vanishes on $\mathcal{M}$ after training. Moreover, we define $F(\mathbf{u}_{i}):=\left\|\mathbf{u}_{i}-D(E(\mathbf{u}_{i}))\right\|_{2}$ instead of $\left\|\mathbf{u}_{i}-D(E(\mathbf{u}_{i}))\right\|_{2}^{2}$ since the former has an $O(1)$ gradient even near $\mathcal{M}$ while the latter’s gradient is of order $o(1)$ near $\mathcal{M}$ which may cause the regularization to be ineffective. The overall regularized loss function of our algorithm in general cases reads
(21) $\min_{\theta}l(\theta):=\min_{\phi_{\theta}}\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\mathbf{y}_{i}-\phi_{\theta}(\mathbf{u}_{i})\right\|_{2}^{2}+\lambda\left(\left(\nabla F(\mathbf{u}_{i})\right)^{T}L(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})\right)^{2}\right).$
@@
In the second term, $D,E$ are optimized in advance and frozen during the training of $\phi_{\theta}$. Since the loss function is regularized by a term that approximately measures the deviation of the velocity vector $L(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})$ from the tangent space of the data manifold, we name our algorithm “MLHS with tangent-space regularized estimator”. The overall algorithm is summarized in Algorithm 1.
There is a subtle difference in our algorithm between linear and nonlinear problems. While in linear case we can directly construct the projection operators onto the orthogonal complement of the linear manifold denoted by $\operatorname{P}_{V^{\perp}}$, in nonlinear case each function $F(\mathbf{u})$ only contributes one vector normal to $\mathcal{M}$. A technique to obtain more normal vectors in the nonlinear case is to have several $F_{i}$ composed by different autoencoders trained with different architectures or initializations. In this case, the regularization term will have the form
(22) $\lambda\left\|\left(\nabla\mathbf{F}(\mathbf{u}_{i})\right)^{T}L(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})\right\|_{2}^{2}=\lambda\sum_{j=1}^{K}\left(\left(\nabla F_{j}(\mathbf{u}_{i})\right)^{T}L(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})\right)^{2}.$
Here $K$ is number of normal directions chosen to be penalized with a maximum value $K_{\max}=m-l$. Notice that a greater $K$ also implies a greater computational and memory cost. Throughout our experiment, we will only use one autoencoder to construct the regularization term as this already provides a significant improvement in the simulation accuracy.